Give the polar coordinates: (4, 0).
Answer by stanbon(48531)

Give the polar coordinates: (4, 0).
r = sqrt(x^2+y^2) = 4
theta = tan^-1(y/x) = tan^-1(0) = 0 degrees
=============
polar coordinates (4,0)
===========================
Cheers,
Stan H.